OpenStudy (emily_98):

Which polynomial is prime? x3 + 3x2 + 2x + 6 x3 + 3x2 – 2x – 6 10x2 – 4x + 3x + 6 10x2 – 10x + 6x – 6 please help

OpenStudy (anonymous):

\[x^3+3x^2+2x+6=x^2(x+3)+2(x+3)=?\] \[x^3+3x^2-2x-6=x^2(x+3)-2(x+3)=?\] \[10x^2-4x+3x+6=10x^2-x+6 \] Factors can not be made. \[10 x^2-10x+6x-6=10 x(x-1)+6(x-1)=?\] can you answer now?

OpenStudy (emily_98):

the third one?

OpenStudy (reemii):

I think so too, given the work done by surjithayer

OpenStudy (anonymous):

correct.